# Perceptions of Linguistic Uncertainty by Language Models and Humans

### v1

@@ 4 Methodology
Our full set of experiments includes experiments with both humans and LLMs as participants and with both non-verifiable (NV) and verifiable (V) statements. The baseline experiment, as described in Section 3, consists of human participants and non-verifiable statements (denoted human + NV). In this section, we describe extensions to include verifiable statements and LLM participants, resulting in three additional sets of experiments: human + V, LLM + V, LLM + NV. To help us draw comparisons between the three additional settings and the baseline experiment, we conclude this section with the description of two quality metrics.
@@ 4.1 Verifiable Statements
In addition to the non-verifiable statements described in Section 3, our dataset also includes verifiable statements, for the purpose of assessing the effects of prior knowledge on quantifying linguistic uncertainty. To this end, we create 60 verifiable statements based on a multiple-choice question-answering dataset from The Question Company.555 https://www.thequestionco.com/ Starting with 30 of the dataset’s “easy” questions and corresponding multiple-choice options, we write true statements that use the correct answer and false statements using one of the incorrect answers. Examples of verifiable statements and additional details about the dataset are included in Appendix B. We analyze results for both humans and LLMs with these 60 verifiable statements in the main paper: for additional validation, our dataset includes 400 additional statements (generated from the AI2-ARC question answering dataset (Clark et al., 2018) via a similar procedure), which we analyze in Appendix E.
@@ 4.2 Numerical Responses from LLMs
To obtain uncertainty estimates from LLMs, we create prompts similar to the queries provided to humans (see list of prompts in Appendix B.4). Our goal is to estimate an empirical distribution, per uncertainty expression $u$, over the LLM’s generated responses, in a manner similar to how empirical distributions for humans are generated (e.g., see Figure 4). In the results in this paper we focus on greedy decoding, where we select the numerical response that has the highest probability in the next-token probability distribution generated by the LLM conditioned on the prompt, i.e., decoding with temperature=0. Because this sampling approach requires no knowledge about the weights or next-token probabilities, it is applicable to any model, including those behind black-box APIs, such as Gemini (Anil et al., 2024) and GPT-4 (Achiam et al., 2024). We focus primarily on this greedy sampling approach since it aligns more closely with the human responses, i.e., we want to assess the ability of each LLM as if it were an individual human providing a single response rather than asking it to match a population distribution of human responses [^6] In Section 5 we evaluate the sensitivity of our conclusions by also analyzing decoding with temperature=1.. We obtain an empirical distribution of LLM responses, for each uncertainty expression $u$, by generating responses in this manner over multiple statements $e$ and speakers $s$. We refer the interested reader to Appendix C for additional information on the extraction methodologies used.
@@ 4.3 Metrics
We treat the empirical distribution obtained for the non-verifiable statements with human participants (described in Section 3) as our reference distribution for evaluation purposes, since it reflects human perceptions of uncertainty expressions in a setting that is designed to be free of prior information or biases about the corresponding statements. For every uncertainty expression $u\in\mathcal{U}$, we define a reference conditional probability distribution $P(k|u)$, $k=0,5,10,\ldots,95,100$. where $P(k|u)$ is the empirical distribution from the baseline experiment. Given a response from any agent, human or LLM, in the context of a particular uncertainty expression $u$ we measure the quality of the response using the reference distribution $P(k|u)$.
The primary quality metric that we propose is Proportional Agreement (PA). PA can be defined as follows: if an agent’s response matches bin $k$ for uncertainty expression $u$, then the PA value for that response is defined as $P(k|u)$, where $P$ is the reference (population) distribution defined above. Intuitively, for an expression $u$, this PA score $P(k|u)$ represents the probability that the agent’s response $k$ agrees with that of a randomly selected individual, and is upper bounded for any expression by $\arg\max_{k}P(k|u)$, i.e., by the mode of the $P(k|u)$. The higher the PA value, the better the quality of the response in terms of agreement with the aggregate human population (as reflected by $P(k|u)$). To compute a single score for a particular LLM or individual human, we average the PA score over multiple responses and over the 14 uncertainty expressions [^7] Note that the PA metric is similar to the log-probability metric widely used to score probabilistic models in machine learning. However, it is not a likelihood in the sense that a likelihood corresponds to measuring the probability mass a model assigns to an observed outcome. Thus, in this non-likelihood context it is appropriate to average the PA scores directly (rather than taking products of probabilities as would be done under an IID likelihood assumption)..
An alternative to the PA metric would be to compare histograms of responses, e.g., based on multiple responses from agents for a particular uncertainty expression $u$. We provide numerical results for histogram comparisons (using the Wasserstein distance between histograms) in Appendix D, but we consider this to be of secondary interest since we are not requiring any LLM or individual human to necessarily replicate the full population variability of responses.
@@ 5 Results
This section examines the ability of several well-known LLMs to interpret uncertainty expressions [^8] We focus in this section on a subset of popular models; results for all 10 models evaluated are in Appendix D.. We begin by assessing models’ abilities to produce numerical responses that resemble human-like trends (e.g., higher numerical responses assigned to higher certainty expressions and vice-versa). We then study the effect of prior knowledge in the perception of uncertainty of both humans and models. We conclude with an assessment of the generalizability of our findings.
@@ 5.1 How well do LLMs perceive uncertainty?
As established in prior work (Wallsten et al., 1986b; Fagen-Ulmschneider, 2019; Willems et al., 2019) and in our baseline experiment (Section 3), humans show population-level agreement in mapping uncertainty expressions to numerical responses. In this section, we assess whether LLMs possess a similar ability to ascribe numerical responses to uncertainty expressions. To this end, we prompt LLMs to provide responses for the same non-verifiable (NV) statements as in the baseline experiment. Figure 5 shows the expression-wise histograms for these responses for two LLMs: GPT-4o and OLMo (7B), each of which can be directly compared to the histogram for humans in Figure 4.
(a) GPT-4o
(b) OLMo-7B Instruct
@@
Figure 5: Model empirical distributions of numerical responses per uncertainty expression in the non-verifiable setting (LLM + NV). Highlighted boxes represent the mode value for each expression. Even though we found no evidence of explicit instruction tuning datasets focusing on uncertainty estimation tasks, our results show that GPT-4o generally manifests human-like behavior, whereas OLMo (7B) does not.
@@
Visually, we observe that GPT-4o matches the human distributions well, with smaller variance per distribution, while OLMo (7B) is less well-matched. In Figure 13 in Appendix D, we see that most LLMs map uncertainty expressions to numerical responses in a manner consistent with human behavior, with higher values for expressions that are perceived by humans as higher-certainty (e.g., “almost certain,” “highly likely”) and lower values for lower-certainty expressions (e.g., “very unlikely”). Only two of the LLMs evaluated, OLMo (7B) and Gemma (2B), fail to reproduce this “increasing” pattern across expressions (relative to the human pattern). One clear difference between humans and LLMs is that the conditional distributions of LLMs have lower entropy (or variance) relative to the human distributions, with the LLM distributions tending to be much more concentrated on a small number of responses compared to the variance in responses from a population of humans.
@@
These observations are reflected more precisely by the PA scores in Table 1. We observe that larger and newer LLMs (in particular, GPT-4, LLama3 (70B), and Gemini) perform especially well on this task under the PA metric, being at 85% or above in terms of matching the modal scores that a human population assigns to each uncertainty expression. In fact, 7 out of the 10 LLMs evaluated are significantly better matched to population modal responses than are individual humans on average [^9] The average performance of individual humans is represented by the Human Individual row in Table 1.. This aligns with the high-level findings of Maloney et al. (2024), in particular, that the difference between the numerical responses of GPT-4 and humans were similar to (or smaller than) inter-human differences. In the context of our experiments, these high scores reflect that LLMs tend to be more consistent than individual humans in terms of agreement with aggregate human responses.
@@
The MAE scores in Table 1, where lower is better, are highly anti-correlated with the PA scores, and tell a similar story in terms of which models perform better. To provide a sense of scale, the MAE numbers are lower-bounded by 0 and upper-bounded by 25 (the expected MAE for random responses).
PA (95% CI) % PA MAE (95% CI)
@@
GPT-4 24.4 (24.2, 24.6) 88.4 4.64 (4.53, 4.74)
GPT-4o 18.9 (18.4, 19.5) 68.5 5.58 (5.34, 5.79)
Gemini 25.4 (25.1, 25.6) 92.0 4.09 (3.92, 4.23)
Llama3 (8B) 17.8 (17.4, 18.2) 64.5 11.99 (10.92, 13.13)
Llama3 (70B) 23.6 (23.4, 23.8) 85.5 5.56 (5.34, 5.80)
Mixtral 8x7B 21.8 (21.4, 22.0) 79.0 5.88 (5.44, 6.34)
Mixtral 8x22B 21.8 (21.6, 22.0) 79.0 7.20 (6.56, 7.81)
OLMo (7B) 11.1 (10.7, 11.5) 40.2 21.41 (20.17, 22.41)
Gemma (2B) 8.1 (7.7, 8.5) 29.3 20.17 (19.11, 21.17)
@@ 5.2 Does knowledge affect uncertainty perceptions of LLMs?
In this section, we assess the extent to which LLMs, and humans, are biased by their prior knowledge or beliefs in mapping uncertainty expressions to numerical responses. To investigate this question we collect responses from humans and LLMs on our verifiable (V) dataset, which includes both true and false common-knowledge statements (based on correct or incorrect answers, respectively, to multiple-choice questions). We find that average PA scores for both humans and LLMs are systematically lower for verifiable statements compared to the non-verifiable responses (Table 2). This suggests that prior knowledge about a statement makes it more difficult to quantify the beliefs of someone else about that statement. While humans show a small drop in their PA score, this reduction in PA is particularly pronounced for LLMs: all 10 LLMs demonstrated a significant reduction in PA, averaging a 4.3 point drop in score (across all models), compared to a 0.9 point drop for humans.
PA (95% CI) $\Delta$ PA MAE (95% CI) $\Delta$ MAE
@@
GPT-4 22.1 (21.7, 22.4) -2.3 3.84 (3.03, 4.45) -0.80
GPT-4o 15.2 (14.5, 15.9) -3.7 7.05 (6.45, 7.62) 1.47
Gemini 21.3 (20.7, 21.8) -4.1 7.23 (6.04, 8.49) 3.14
Llama3 (8B) 10.1 (9.5, 10.7) -7.7 16.59 (15.10, 18.18) 4.60
Llama3 (70B) 18.9 (18.1, 19.6) -4.7 13.73 (11.92, 15.64) 8.17
Mixtral 8x7B 15.2 (14.5, 15.9) -6.6 12.23 (10.37, 14.18) 6.35
Mixtral 8x22B 18.6 (18.3, 19.0) -3.2 9.78 (8.51, 11.08) 2.58
OLMo (7B) 7.6 (7.2, 8.0) -3.5 33.66 (31.82, 35.11) 12.25
Gemma (2B) 5.3 (5.0, 5.6) -2.8 25.07 (23.45, 26.66) 4.9
To investigate these differences in more detail, we consider the mean response values produced by the 6 models exhibiting the highest PA score. These values differ systematically depending on whether the statement is true or false: across the 6 LLMs in Figure 6, the mean response is 7.0 percentage points lower for false than true statements. This indicates that the models assign higher response values to the same uncertainty expression when they believe the associated statement it refers to is true than when they believe it to be false (providing a quantitative validation of the ChatGPT example in Figure 1 in Section 1).
Figure 6: Mean numerical response for the verifiable statements discriminated by truthfulness of statements. The mean numerical responses produced by LLMs when evaluated in the context of true statements is significantly larger than when evaluated in the context of false statements. This difference is much larger in magnitude than the difference shown by a human population.
@@
(c) “unlikely”
Figure 7: Mean response for verifiable statements discriminated by truthfulness of statements for three different uncertainty expressions. We observe that the differences in mean numerical responses differs by uncertainty expression. Despite changes in magnitude, ChatGPT systematically exhibits the largest gap.
Overall, we find that all of the evaluated LLMs demonstrate significant biases based on their prior knowledge, well beyond those of humans. Our results indicate that when an LLM believes a statement is false, it tends to perceive the speaker’s certainty as low, regardless of the actual uncertainty expressed by the speaker (and vice versa), lacking “theory of mind." Our findings in general are also consistent with recent results in the context of LLM reading comprehension, where the efficacy of LLM comprehension is sensitive to the degree of prior knowledge it has about the text it is analyzing Basmov et al. (2024).

### v2

@@ 4 Methodology
Our full set of experiments includes experiments with both humans and LLMs as participants and with both non-verifiable (NV) and verifiable (V) statements. The baseline experiment, as described in Section 3, consists of human participants and non-verifiable statements (denoted human+NV). In this section, we describe extensions to include verifiable statements and LLM participants, resulting in three additional sets of experiments: human+V, LLM+V, LLM+NV. To help us draw comparisons between the three additional settings and the baseline experiment, we conclude this section with the description of two metrics.
@@ 4.1 Verifiable Statements
In addition to the non-verifiable statements described in Section 3, our dataset also includes verifiable statements, for the purpose of assessing the effects of prior knowledge on quantifying linguistic uncertainty. To increase the chances that both LLMs and humans are familiar with the verifiable statements, we focus on concise, general-knowledge statements based on widely recognized facts (e.g., geography, history of art, science). Specifically, we create 60 verifiable statements based on a multiple-choice question-answering trivia dataset from The Question Company.555 https://www.thequestionco.com/ Starting with 30 of the dataset’s “easy” questions and corresponding multiple-choice options, we write true statements that use the correct answer and false statements using one of the incorrect answers. Examples of verifiable statements and additional details about the dataset are included in Appendix B. We focus on results for both humans and LLMs with these 60 verifiable statements in the main paper but include a validation analysis using 400 additional statements in Section 5.3. These additional statements are extracted via a similar procedure from AI2-ARC (Clark et al., 2018), a grade-school level, multiple-choice science question answering dataset used to evaluate state-of-the-art LLMs’ reasoning capabilities (Beeching et al., 2023; Jiang et al., 2023; Achiam et al., 2024).
@@ 4.2 Numerical responses from LLMs
To obtain uncertainty estimates from LLMs, we create prompts similar to the queries provided to humans (see Appendix C). Our goal is to estimate an empirical distribution, per uncertainty expression $u$, over each LLM’s generated responses, in a manner similar to how empirical distributions for humans are generated (e.g., see Figure 3). In the results in this paper we focus on greedy decoding, where we select the numerical response that has the highest probability in the next-token probability distribution generated by the LLM conditioned on the prompt, i.e., decoding with temperature=0. Because this sampling approach requires no knowledge about the weights or next-token probabilities, it is applicable to any model, including those behind black-box APIs, such as Gemini (Anil et al., 2024) and GPT-4 (Achiam et al., 2024). While focusing primarily on this greedy sampling approach allow us to efficiently compare the modal behavior of different LLM model families in equal terms (regardless of the available information), it does not provide insight into distributional behavior. Thus, in Section 5 we include results using probabilistic decoding with temperature=1 to evaluate the sensitivity of our conclusions to decoding method. Additional information on the extraction methodologies used can be found in Appendix D.
@@ 4.3 Metrics
We treat the empirical distribution obtained for the non-verifiable statements with human participants (described in Section 3) as our reference distribution for evaluation purposes, since it reflects human perceptions of uncertainty expressions in a setting that is designed to be free of prior information or biases about the corresponding statements. For every uncertainty expression $u\in\mathcal{U}$, we define a reference conditional probability distribution $P(k|u)$, $k=0,5,10,\ldots,95,100$, where $P(k|u)$ is the empirical distribution from the baseline experiment. Given a response from any agent, human or LLM, in the context of a particular uncertainty expression $u$ we measure the quality of the response using the reference distribution $P(k|u)$.
The primary quality metric that we propose is Proportional Agreement (PA). PA can be defined as follows: if an agent’s response matches bin $k$ for uncertainty expression $u$, then the PA value for that response is defined as $P(k|u)$, where $P$ is the reference (population) distribution defined above. Intuitively, for an expression $u$, this PA score $P(k|u)$ represents the probability that the agent’s response $k$ agrees with that of a randomly selected individual, and is upper bounded for any expression by $\arg\max_{k}P(k|u)$, i.e., by the mode of the $P(k|u)$ values. The higher the PA value, the better the quality of the response in terms of agreement with the aggregate human population (as reflected by $P(k|u)$). To compute a single score for a particular LLM or individual human, we average the PA score over multiple responses and over the 14 uncertainty expressions. [^6] Note that the PA metric is similar to the log-probability metric widely used to score probabilistic models in machine learning. However, it is not a likelihood in the sense that a likelihood corresponds to measuring the probability mass a model assigns to an observed outcome. Thus, in this non-likelihood context it is appropriate to average the PA scores directly (rather than taking products of probabilities as would be done under an IID likelihood assumption).
One drawback of using the PA metric is that it penalizes deviations from the reference distribution’s mode. As a result, it may fail to capture the nuanced distributional differences that emerge when conveying uncertainty about events with varying severity or base rate (Wallsten et al., 1986a; Weber and Hilton, 1990; Willems et al., 20). However, in our experiments, we do not expect to observe systematic distributional differences, since the uncertainty expressions are used in neutral contexts that concern unknown people and extraneous events/facts. Nonetheless, we include an alternative to the PA metric that compares histograms of responses, e.g., based on multiple responses from agents for a particular uncertainty expression $u$. In particular, we provide numerical results for histogram comparisons (using the Wasserstein distance between histograms) in Appendix E.
@@ 5 Results
This section examines the ability of several well-known LLMs to interpret uncertainty expressions. We begin by assessing models’ abilities to produce numerical responses that resemble human-like trends (e.g., higher numerical responses assigned to higher-certainty expressions and vice-versa). We then study the effect of prior knowledge in the perception of uncertainty of both humans and models. We conclude with an assessment of the generalizability of our findings.
@@ 5.1 How well do LLMs perceive uncertainty?
As established in prior work (Wallsten et al., 1986b; Fagen-Ulmschneider, 2019; Willems et al., 20) and in our baseline experiment (Section 3), humans show population-level agreement in mapping uncertainty expressions to numerical responses. In this section, we assess whether LLMs possess a similar ability to ascribe numerical responses to uncertainty expressions. To this end, we prompt LLMs to provide responses for the same non-verifiable (NV) statements as in the baseline experiment. Figure 4 shows the expression-wise histograms for these responses for two LLMs: GPT-4o and OLMo (7B), each of which can be directly compared to the histogram for humans in Figure 3.
@@
Figure 4: Model empirical distributions of numerical responses per uncertainty expression in the non-verifiable setting (LLM+NV). Highlighted boxes represent the mode value for each expression. Even though we found no evidence of explicit instruction tuning datasets focusing on uncertainty estimation tasks, these results suggest that GPT-4o generally manifests human-like behavior, whereas OLMo (7B) does not.
@@
Visually, we observe that GPT-4o matches the human distributions well, with smaller variance per distribution, while OLMo (7B) is less aligned. In Figure 14 in Appendix E, we show that most LLMs map uncertainty expressions to numerical responses in a manner consistent with human behavior, with higher values for expressions that are perceived by humans as higher-certainty (e.g., “almost certain,” “highly likely”) and lower values for lower-certainty expressions (e.g., “very unlikely”). Only two of the LLMs evaluated, OLMo (7B) and Gemma (2B), fail to reproduce this “increasing” pattern across expressions. One clear difference between humans and LLMs is that the conditional distributions of LLMs have lower entropy (or variance) relative to the human distributions, with the LLM distributions tending to be much more concentrated [^7] Appendix E.3 shows evidence that greedy decoding results in lower variability in estimated distributions compared to humans. Additionally, while probabilistic decoding (temperature=1) generally increases variability for most models, this effect is not observed for GPT-4. on a small number of responses compared to the variance in responses from a population of humans.
@@
These observations are reflected more precisely by the PA scores in Tables 1 and 9. We observe that larger and newer LLMs (in particular, GPT-4, LLama3 (70B), and Gemini) perform especially well on this task under the PA metric, being at 85% or above in terms of matching the modal scores that a human population assigns to each uncertainty expression. In fact, 7 out of the 10 LLMs evaluated are significantly better matched to population modal responses than are individual humans on average [^8] The average performance of individual humans is represented by the Human Individual row in Table 1.. This aligns with the high-level findings of Maloney et al. (2024), in particular, that the difference between the numerical responses of GPT-4 and humans were similar to (or smaller than) inter-human differences. In the context of our experiments, these high scores reflect that LLMs tend to be more consistent than individual humans in terms of agreement with aggregate human responses.
@@
The MAE scores in Table 1 (lower is better) are highly anti-correlated with the PA scores and tell a similar story in terms of which models perform better. To provide a sense of scale, the MAE numbers are lower-bounded by 0 and upper-bounded by 25 (the expected MAE for random responses).
Table 1: Human-LLM agreement for non-verifiable statements. Average Proportional Agreement (PA), PA as a fraction of the Human Mode results (PA), and absolute error between mean responses (MAE). Human Mode represents the mode of the human NV distribution, whereas Human Individual represents the PA score of individual human responses relative to the population.
PA % PA MAE
@@
GPT-4 $24.4$ 88.4 $4.64$
GPT-4o $18.9$ 68.5 $5.58$
Gemini $25.4$ 92.0 $4.09$
LLama3 (8B) $17.8$ 64.5 11.9911.99{11.99}11.99
LLama3 (70B) $23.6$ 85.5 5.565.56{5.56}5.56
Mixtral 8x7B $21.8$ 79.0 5.885.88{5.88}5.88
Mixtral 8x22B $21.8$ 79.0 7.207.20{7.20}7.20
OLMo (7B) $11.1$ 40.2 21.4121.41{21.41}21.41
Gemma (2B) $8.1$ 29.3 20.1720.17{20.17}20.17
@@ 5.2 Does knowledge affect uncertainty perceptions of LLMs?
In this section, we assess the extent to which LLMs, and humans, are biased by their prior knowledge or beliefs in mapping uncertainty expressions to numerical responses. To investigate this question we collect responses from humans and LLMs on our verifiable (V) dataset, which includes both true and false common-knowledge statements (based on correct or incorrect answers, respectively, to multiple-choice questions). We find that average PA scores for both humans and LLMs are systematically lower for verifiable statements compared to the non-verifiable responses (Tables 2 and 10). This suggests that prior knowledge about a statement [^9] We validempirically ated our use of true correctness as a proxy for the LLMs’ beliefs in Appendix B.3. makes it more difficult to quantify the beliefs of someone else about that statement. While humans show a small drop in their PA score, this reduction in PA is particularly pronounced for LLMs: all 10 LLMs demonstrated a significant reduction in PA, averaging a 4.3 point drop in score (across all models), compared to a 0.9 point drop for humans.
Table 2: Human-LLM agreement for verifiable statements. Average Proportional Agreement (PA), absolute error between mean responses (MAE), and the difference between these scores and those from the non-verifiable statements (Table 1) ($\Delta$ PA and $\Delta$ MAE, respectively). Again, Human Mode represents the mode of the human NV distribution, whereas Human Individual represents the average behavior across individual humans on the verifiable setting.
PA $\Delta$ PA MAE $\Delta$ MAE
@@
GPT-4 ${22.1}$2.3 ${3.84}$0.80
GPT-4o ${15.2}$3.7 7.057.05{7.05}7.05 1.47
Gemini ${21.3}$4.1 7.237.23{7.23}7.23 3.14
LLama3 (8B) ${10.1}$7.7 16.5916.59{16.59}16.59 4.60
LLama3 (70B) ${18.9}$4.7 13.7313.73{13.73}13.73 8.17
Mixtral 8x7B ${15.2}$6.6 12.2312.23{12.23}12.23 6.35
Mixtral 8x22B ${18.6}$3.2 9.789.78{9.78}9.78 2.58
OLMo (7B) ${7.6}$3.5 33.6633.66{33.66}33.66 12.25
Gemma (2B) ${5.3}$2.8 25.0725.07{25.07}25.07 4.9
To investigate these differences in more detail, we consider the mean response values produced by the 6 models exhibiting the highest PA score. These values differ systematically depending on whether the statement is true or false: across the 6 LLMs in Figure 5, the mean response is 7.0 points lower for false than true statements. This indicates that the models assign higher response values to the same uncertainty expression when they believe the associated statement is true than when they believe it to be false (providing a quantitative validation of the ChatGPT example in Section 1).
Figure 5: Mean numerical response for the verifiable statements discriminated by truthfulness of statements. The mean numerical responses produced by LLMs when evaluated in the context of true statements is significantly larger than when evaluated with the false statements. This difference is much larger in magnitude than the difference shown by a human population.
@@
Figure 6: Mean response for verifiable statements discriminated by truthfulness of statements for three different uncertainty expressions. We observe that the differences in mean numerical responses differ by uncertainty expression. Despite changes in magnitude, ChatGPT systematically exhibits the largest gap.
Overall, we find that all of the evaluated LLMs demonstrate significant biases based on their prior knowledge, well beyond those of humans. Our results indicate that when an LLM believes a statement is false, it tends to perceive the speaker’s certainty as low, regardless of the actual uncertainty expressed by the speaker (and vice versa), lacking “theory of mind.” Our findings in general are also consistent with recent results in the context of LLM reading comprehension, where the efficacy of LLM comprehension is sensitive to the degree of prior knowledge it has about the text it is analyzing Basmov et al. (2024).